Bet on a single space at roulette until you win, then stop.

What is the chance you will play a dozen times?

What is the chance you will play no more than a dozen times?
Answer #1

Probability of winning on a single play =1/38

1)

chance you will play a dozen times =P(lose in first 11 games and won on 12 th turn)

=(1-1/38)11*(1/38) =0.0196

2)

P( play no more than a dozen times ) =1-P(no win in first 12 games) =1-(37/38)12 =0.2739